# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ Appendix C Supporting Lemmas
$\leq 2\sqrt{\lambda_{j}}\cdot\|{\bm{\xi}}_{j}\widehat{\bm{\Gamma}}_{b}\|\cdot\|{\bm{\Delta}}_{bb}\|_{{\mathop{\mathrm{op}}}}\leq\frac{4\sqrt{2c}}{\lambda_{\min}(\mathbf{Q})}$
almost surely for all $n$ large enough. We hence conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, $|f^{\prime}_{n}(z)|$ is uniformly bounded. Combining with the case $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, we have that $|f^{\prime}_{n}(z)|$ is uniformly bounded when $z\geq 0$. After similar argument in the proof of Lemma C.6 completes the proof of Lemma C.7.
@@
We then give the proof of Theorem D.1. In the proof of Theorem D.1, we give the precise asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, and then construct statistics only related to $\mathbf{X}$ which share the same asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, finally we use these statistics to estimate $SR(\mathbf{Q})$. Define
@@
Recall the definition of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$ in (2.2) and (2.3) that
@@
The next proposition further indicates us how to construct the statistics only related to $\mathbf{X}$ which share the same asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$.
@@
Proposition D.3 constructs the estimation of the key part in the asymptotic of $T_{n,2}(\mathbf{Q})$ given in Proposition D.2, and is proved in Appendix D.2. Now, the only thing remained for us is to construct statistics which give precise estimations of $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ respectively. The next proposition established such precise estimations.
@@
Below we prove the three conclusions one by one.
The first conclusion. If $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded, from Proposition D.2 and D.3, we conclude that
@@
$\leq\frac{C_{\mathbf{Q}}(1+s_{0})}{C_{\mathbf{Q}}(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}.$ (D.6)
Here, $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ denote by the largest and smallest eigenvalue of the matrix. The first inequality is by ${\mathrm{tr}}(\mathbf{A}\mathbf{B})\leq\lambda_{\max}(\mathbf{A}){\mathrm{tr}}\mathbf{B}$ for $\mathbf{A},\mathbf{B}\geq 0$, the second inequality is by $\lambda_{\max}\Big{(}\mathbf{Q}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\Big{)}=\lambda_{\max}\Big{(}\Big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}^{-1}\cdot{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}$. Therefore, plugging (D.6) into (D.5), we have
$g({\bm{\Sigma}})$ $\geq\frac{1+s_{0}}{C_{\mathbf{Q}}(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
@@
$\frac{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}{\sqrt{g({\bm{\Sigma}})}}$ $\leq\sqrt{\frac{C_{\mathbf{Q}}(1+s_{0})+1}{1+s_{0}}}\cdot\sqrt{{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}}$
@@
Therefore we conclude that $SR(\mathbf{Q})$ is almost surely bounded, which completes the proof of the third conclusion in Theorem D.1. ∎
@@
Direct calculation gives us
@@
$\qquad\qquad=-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A},$
@@
$\bigg{|}\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}-\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}\bigg{|}$
$\quad=\bigg{|}T_{n,2}(\lambda)-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}\bigg{|}$
$\quad~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
@@
$s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},\quad s_{1,{\bm{\Sigma}}}=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},$
@@
$\qquad=\lambda_{\min}\bigg{(}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{)}\geq\frac{1+s_{0}}{C_{\mathbf{Q}}(1+s_{0})+1}.$
@@
the conclusion directly holds from the fact that the maximum and minimum eigenvalue of $\big{(}\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}$ can be upper and lower bounded by a constant, respectively. This completes the proof of Lemma E.2.
@@
Appendix F Relationship Between Ridge Regularization and Maximum Sharpe Ratio
According to the theory proposed by Markowitz (1952), the maximum population Sharpe ratio can be expressed as:
@@ Appendix F Relationship Between Ridge Regularization and Maximum Sharpe Ratio
The proof of Proposition F.1 is given in Section F.1. This proposition tells us that a well-designed regularization matrix $\widetilde{\mathbf{Q}}$ allows $SR(\widetilde{\mathbf{Q}})$ to closely approximate the maximum population Sharpe ratio. Furthermore, we could use the criteria established in Theorem 2.7 to select $\mathbf{Q}$ that approximately maximizes $SR(\mathbf{Q})$.
Another interesting case arises when ${\bm{\Sigma}}$ exhibits a specific factor structure, defined as ${\bm{\Sigma}}=\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}$. Here, $\mathbf{B}=(\bm{b}_{1},\dots,\bm{b}_{K})\in\mathbb{R}^{p\times K}$ with each $\|\bm{b}_{j}\|_{2}=\Theta(\sqrt{p})$ for $j\in[K]$, $\operatorname{rank}(\mathbf{B})=K$ and $\mathbf{D}$ is a positive matrix with eigenvalues bounded away from 00 and +$\infty$+ ∞. Under this model, the following proposition provides further insight into the role of $\mathbf{Q}$.
@@
The proof of Proposition F.2 is given in Section F.2. This proposition demonstrates that even with the presence of factors, i.e. spiked eigenvalues, it is still possible to identify a regularization matrix $\widetilde{\mathbf{Q}}$ with a bounded operator norm that can approximate the maximum population Sharpe ratio. Again, we can utilize Theorem 2.7 to evaluate the effectiveness of the design of $\mathbf{Q}$.
@@
$\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\cdot\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}\cdot\frac{1}{bmu/SR(\mathbf{Q})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
where $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ are defined in Lemma C.5 by letting $z=0$. We first consider the expression of ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ and ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ by specifically choosing $\mathbf{Q}=q\cdot D$. Letting $\lambda=1+(1+s_{0})q$, we have

### v3

@@ Appendix C Supporting Lemmas
$\leq 6C$
almost surely for all $n$ large enough and for some large enough absolute constant $C>0$. We hence conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$, $|f^{\prime}_{n}(z)|$ is uniformly bounded. Hence we have that $|f^{\prime}_{n}(z)|$ is uniformly bounded when $z\geq 0$. After similar argument in the proof of Lemma C.6 completes the proof of Lemma C.7.
@@
In order to prove Theorem D.1, we give the precise asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, and then construct a statistics only using in-sample data which share the same asymptotic as $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, finally we use these statistics to estimate $SR(\mathbf{Q})$. Define
@@
Recall the definitions of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$ in (2.2) and (2.3) that
@@
The next proposition further indicates how to construct an estimator for $g({\bm{\Sigma}})$ only using $\widehat{\bm{\Sigma}}$ instead of ${\bm{\Sigma}}$ to achieve the same asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$.
@@
Proposition D.3 constructs the estimation for the key part in the asymptotic of $T_{n,2}(\mathbf{Q})$ given in Proposition D.2, and is proved in Appendix D.2. Now, the only thing remains for us is to construct estimators for $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ respectively. The next proposition established such precise estimations.
@@
Below we first prove the three conclusions when $\mathbf{Q}\neq{\bm{0}}$.
The first conclusion. If $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded, from Proposition D.2 and D.3, we conclude
@@
$\leq\frac{C(1+s_{0})}{C(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}.$ (D.6)
Here, $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ denote by the largest and smallest eigenvalue of the matrix. $C$ is the maximal eigenvalue of ${\bm{\Sigma}}^{-1/2}\mathbf{Q}{\bm{\Sigma}}^{-1/2}$. The first inequality is by ${\mathrm{tr}}(\mathbf{A}\mathbf{B})\leq\lambda_{\max}(\mathbf{A}){\mathrm{tr}}\mathbf{B}$ for $\mathbf{A},\mathbf{B}\geq 0$, the second inequality is by $\lambda_{\max}\Big{(}\mathbf{Q}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\Big{)}=\lambda_{\max}\Big{(}\Big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}^{-1}\cdot{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}$. Therefore, plugging (D.6) into (D.5), we have
$g({\bm{\Sigma}})$ $\geq\frac{1+s_{0}}{C(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
@@
$\frac{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}{\sqrt{g({\bm{\Sigma}})}}$ $\leq\sqrt{\frac{C(1+s_{0})+1}{1+s_{0}}}\cdot\sqrt{{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}}$
@@
Therefore we conclude that $SR(\mathbf{Q})$ is almost surely bounded, which completes the proof of the third conclusion in Theorem D.1.
As for the case $\mathbf{Q}={\bm{0}}$ with $c<1$. The proofs of Lemma C.5 and C.6 when $\mathbf{Q}={\bm{0}}$ are slightly different. The solution of $s_{0}$ in Lemma C.5 can be obtained directly by the cancellation of ${\bm{\Sigma}}$, and $|f_{n}(z)|$ in Lemma C.6 can also be bounded by $2\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathrm{tr}}}/(1-\sqrt{c})^{2}$. Then when $\mathbf{Q}={\bm{0}}$ under $c<1$, it is clear that $1+s_{0}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1/(1-c)$ and $\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{1+s_{0}+s_{1,\mathbf{Q}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1/(1-c)^{2}$ after simple algebra calculation. Following very similar steps, we can easily prove the same three conclusions. So we omit some details here. Readers may note that
$1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1-c$
as $n$ tends to infinity when $\mathbf{Q}={\bm{0}}$. The general form of $\widehat{T}_{n,2}(\mathbf{Q})$
$\widehat{T}_{n,2}(\mathbf{Q})=\frac{{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}}{\big{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\big{)}^{2}}$
covers the case $\mathbf{Q}={\bm{0}}$. This completes the proof of Theorem 2.7. ∎
@@
Direct calculations give us
@@
$\qquad=-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A},$
@@
$\bigg{|}T_{n,2}(\lambda)-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}\bigg{|}$
$\quad=\bigg{|}\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}-\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}\bigg{|}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
@@
$s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},$
$s_{1,{\bm{\Sigma}}}=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},$
@@
$\qquad=\lambda_{\min}\bigg{(}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{)}\geq\frac{1+s_{0}}{C^{\prime}(1+s_{0})+1}.$
@@
the conclusion directly holds from the fact that the maximal and minimal eigenvalue of $\big{(}\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}$ can be upper and lower bounded by a constant, respectively. This completes the proof of Lemma E.2.
@@
Appendix F Relationship Between Ridge Regularization and Maximal Sharpe Ratio
According to the theory proposed by Markowitz (1952), the maximal population Sharpe ratio can be expressed as:
@@ Appendix F Relationship Between Ridge Regularization and Maximal Sharpe Ratio
The proof of Proposition F.1 is given in Section F.1. This proposition tells us that a well-designed regularization matrix $\widetilde{\mathbf{Q}}$ allows $SR(\widetilde{\mathbf{Q}})$ to closely approximate the maximal population Sharpe ratio. Furthermore, we could use the criteria established in Theorem 2.7 to select $\mathbf{Q}$ that approximately maximizes $SR(\mathbf{Q})$.
Another interesting case arises when ${\bm{\Sigma}}$ exhibits a specific factor structure, defined as ${\bm{\Sigma}}=\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}$. Here, $\mathbf{B}=(\bm{b}_{1},\dots,\bm{b}_{K})\in\mathbb{R}^{p\times K}$ is knonwn as the loading matrix with each $\|\bm{b}_{j}\|_{2}=\Theta(\sqrt{p})$ for $j\in[K]$, $\operatorname{rank}(\mathbf{B})=K$. $\mathbf{D}$ is known as the residual covariance matrix, which is a positive definite matrix with eigenvalues bounded away from 00 and +$\infty$+ ∞. $\mathbf{D}$ is typically assumed to be a diagonal matrix, as the off-diagonal correlations have been taken care of mostly by the low-rank factor component $\mathbf{B}\mathbf{B}^{\top}$. But for the discussion here, we only need to assume $\mathbf{D}$ has bounded spectrum from above and below. Under this model, the following proposition provides further insight into the role of $\mathbf{Q}$.
@@
The proof of Proposition F.2 is given in Section F.2. This proposition demonstrates that even with the presence of factors, i.e. spiked eigenvalues, it is still possible to identify a regularization matrix $\widetilde{\mathbf{Q}}$ with a bounded operator norm that can approximate the maximal population Sharpe ratio. Again, we can utilize Theorem 2.7 to evaluate the effectiveness of the design of $\mathbf{Q}$.
@@
where $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ are defined in Lemma C.5 by letting $z=0$. We first consider the expression of ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ and ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ by specifically choosing $\mathbf{Q}=q\cdot\mathbf{D}$. Letting $\lambda=1+(1+s_{0})q$, we have